One leg of a right traingle is 3 inches shorter than the hypotenuse. The other leg is equal to 8 inches. How long is the first leg?
The correct answer is 55/6 inches, but I can't figure out how ot get it. Please help me!

Answer by stanbon(75887)

One leg of a right traingle is 3 inches shorter than the hypotenuse. The other leg is equal to 8 inches. How long is the first leg?
leg 1 : 8
hypotenuse : x
leg 2 : x-3
----
Pythagoras:
8^2 + (x-3)^2 = x^2
64 + x^2-6x+9 = x^2
6x = 73
x = 73/6 (hypotenuse)
---
x-3 = (73/6)-(18/6) = 55/6
===========
Cheers,
Stan H.